OpenStudy (anonymous):

Solve for x: 2logv4(5x)=3

OpenStudy (anonymous):

2 log 4 (5x) = 3 log 4 (5x) = (3/2) 4^ (log 4(5x)) = 4^(3/2) 5x = 8 x = 8/5

OpenStudy (anonymous):

thank you so much!